I am confused, can I get some help with these please?
Solve each irrational equation. If an equation has no solution, so state:
a. 3/x - 1/6 = 1/x

b. 13/y-1 - 3 = 1/y-1

Answer by Nate(3500)

a.) Remember: anything divided by zero is undefined....x can not be zero
3/x - 1/6 = 1/x multiply all the terms by x
3 - x/6 = 1 multiply all the terms by 6
18 - x = 6
-x = -12
x = 12
b.) y can not equal positive one
13/(y - 1) - 3 = 1/(y - 1)
13 - 3(y - 1) = 1
13 - 3y + 3 = 1
-3y = -15
y = 5